# Dingyi Zhao1, 7, Keming Wu6, Haozhe Wang8, Ping Nie9, Yan Teng1, 🖂, Yingchun Wang1

### v1

@@ 4.3 Retrieval Trustworthiness
Retrieval Trustworthiness evaluates the credibility of external information retrieval and usage in response reports. Modeling based on the hit rate of TSLs and using confidence enhancement mechanism inspired by multiplicative fusion in Bayesian updating, this approach transforms match rates into multiplicative scoring factors to increase the evaluative weight of citation quality. Specifically, matches is categorized into full matches and hostname matches, where $\text{Rate}_{\text{full\,hit}}$ serves as a recall-like metric capturing the precise coverage of provided TSLs, and $\text{Rate}_{\text{host\,hit}}$ reflects the proportion of generalized mentions whose annotation links share the same source domains as the recommended references. These two are assigned different weights and combined to compute the TrustworthyBoost factor as follows:
$\text{TrustworthyBoost}=1+\eta\cdot\bigg[\ \theta\cdot\underbrace{\left(\frac{\text{match}_{\text{full}}}{S}\right)}_{\text{Rate}_{\text{full\,hit}}}+\kappa\cdot\underbrace{\left(\frac{\text{match}_{\text{host}}}{T+1}\right)}_{\text{Rate}_{\text{host\,hit}}}\ \bigg]\in[1,1+\eta]$
where $\text{match}_{\text{full}}$ indicates the number that have been exactly matched in TSLs; $\text{match}_{\text{host}}$ refers to the number of annotation links sharing the same hostname as the TSLs; $S$ and $T$ denote the sizes of the TSLs and the annotations respectively; $\theta$ and $\kappa$ represent the weights for full and hostname matches respectively, with $\theta+\kappa=1$. The coefficient $\eta=0.2$ controls the magnitude of the confidence boost, thereby preventing excessive inflation due to high confidence and avoiding complete nullification when confidence is low. This mechanism preserves scoring stability while enhancing evaluative sensitivity to verifiability and source reliability, thereby improving the responsiveness to external evidence signals.
@@ 4.4 Integrated Scoring Framework
The integrated scoring metric adopts a multiplicative weighting model to enable multidimensional evaluation of report-style generation tasks. The calculation formula is defined as:
$\text{IntegratedScore}=\text{Quality}\times(1-\text{SemanticDrift})\times\text{TrustworthyBoost}\times 100\in[0,120]$
Here, Quality assesses the structural integrity and content quality of the report, SemanticDrift reflects the degree of thematic deviation and is transformed into a positive scoring factor via $1-\text{SemanticDrift}$, and TrustworthyBoost enhances the credibility weight based on authoritative link coverage. This design logic penalizes semantic drift and rewards external support, producing a normalized score on a 100-point scale.
@@
[…]
4.5 Additional Metrics
To evaluate cost-effectiveness under limited resources, the Contribution/Token metric measures model efficiency based on actual token expenditure from reasoning and generation.
5 Experiments
5.1 Overview
We evaluated a total of thirteen models under Rigorous Benchmark, including five DRAs (o3-deep-research-2025-06-26, qwen-deep-research, sonar-deep- research, grok-4-0709-search, and o4-mini-deep-research-2025-06-26), one advanced agent (kimi-k2-0905-preview), and seven web-search-tool-enhanced reasoning models (gemini-2.5-pro, gpt-5-2025-08-07, gpt-4o-search-preview-2025- 03-11, gpt-4.1-2025-04-14, claude-opus-4-1-20250805, claude-sonnet-4- 20250514, and claude-3-7-sonnet-20250219).
To eliminate the randomness, all models with adjustable temperature were evaluated at $\text{temperature}=0.0$. For non-DRA models without embedded annotations, reports were merged with annotations during evaluation. gpt-4o-2024-11-20 independently scored each rubric to evaluate semantic quality. Topical focus was assessed on pure report text without annotations to avoid interference from annotation titles. Related prompts are provided in Appendix A.4. Retrieval Trustworthiness was computed from pure annotations, excluding parameters, anchors, and duplicates to ensure fair matching. Manual verification was randomly conducted on approximately 35% of the scores judged by LLMs, yielding a 99.3% agreement with human evaluations.
5.2 Leaderboard
Table 1: Results with the highest score in bold and the second-highest underlined per column.
Models Quality 11-SDrift TBoost InteScore Usage C/Token
@@ 5.2 Leaderboard
Table 1 and Appendix A.5 presents the leaderboard results, with models ranked in descending order of IntegratedScore. In terms of IntegratedScore, Qwen ranked first, demonstrating strong performance across all dimensions. Sonar followed closely, achieving the highest score in topical focus. Notably, Kimi-K2, a Mixture-of-Experts architecture agent with 1T parameters, attained the highest score in the quality, outperforming all DRAs. For topical focus, Sonar, Qwen, and o3 formed the leading cluster. GPT-5 achieved the highest score in citation reliability, indicating strong external support alignment. In terms of resource consumption, o3 and o4-mini averaged 23K and 18K tokens per report respectively, placing them lowest in contribution efficiency.
Overall, DRAs exhibited stable semantic quality and control in structured report generation, while web-search-tool-augmented models showed promise in external support and efficiency. These results validate the Rigorous Bench’s ability to differentiate model capabilities across multiple dimensions, providing a robust empirical foundation for future optimization of DRAs.
@@ 5.3 Supplementary Dimensions
As shown in Table 2, GPT-4.1 showed minimal retrieval activity, with an average of only 0.39 times. In contrast, both o4-mini and o3 displayed intensive inference and retrieval patterns, suggesting more complex reasoning chains and information acquisition strategies. In terms of retrieval efficiency, o3 slightly exceeded o4-mini in total retrieved links and annotations, while o4-mini achieved marginally higher RetrievalIndex, reflecting stronger filtering and citation precision, as shown in Table 3.
Table 2: Average inference and retrieval times.
@@
Table 3: RetrievalIndex of o3 and o4-mini.
Models $\text{{num}}_{\text{{retr}}}$ $\text{{num}}_{\text{{anno}}}$ RIndex
@@ 7 Conclusions
In this paper, we present the Rigorous Bench and the multidimensional evaluation framework that systematically assesses the performance and capabilities of DRAs. By leveraging challenging queries across diverse thematic domains and high-quality reference bundles, our framework enables rigorous evaluation of report-style outputs along axes of semantic quality, topical focus, and retrieval trustworthiness. Empirical results show that contemporary DRAs substantially outperform conventional tool-augmented models in complex task scenarios, while also exposing key limitations and trade-offs. These insights elucidate current challenges in DRA design and provide a foundation for the development of DRAs as efficient, stable, and interpretable intelligent agents.
Appendix A Appendix
A.1 Taxonomy of Domains
@@ 00 Unclassified Cannot be categorized 3
A.2 General-Report Rubrics
A complete listing of the 48 GRRs, used to evaluate general reporting quality, is shown in Figure 4.
@@ A.2 General-Report Rubrics
A.3 Examples of Entries
To illustrate the precision and rigor of our benchmark, we present four representative entries: ID 040216, 07001, 08004, and 10236. These entries span distinct domains and query types and serve as concrete examples of structural completeness, rubric coverage, and citation fidelity.
Figure 5: Example ID 040216 of benchmark entries.
Figure 6: Examples ID 07001 and 10236 of benchmark entries.
@@ A.3 Examples of Entries
A.4 Scoring Prompts
@@ 2. A keyword to evaluate Your task is to: - Read the report carefully
A.5 Supplementary Experimental Observations
A.5.1 Quality Share Proportion
Figure 8: Quality share proportion across models.
@@ A.5.1 Quality Share Proportion
A.5.2 Evaluation Across Domains

### v2

@@ 4.3 Retrieval Trustworthiness
Retrieval Trustworthiness evaluates the credibility of external information retrieval and usage in response reports. Modeling based on the hit rate of TSLs and using confidence enhancement mechanism inspired by multiplicative fusion in Bayesian updating, this approach transforms match rates into multiplicative scoring factors to increase the evaluative weight of citation quality. Specifically, matches are categorized into full matches and hostname matches, where $\text{Rate}_{\text{full\,hit}}$ serves as a recall-like metric capturing the precise coverage of provided TSLs, and $\text{Rate}_{\text{host\,hit}}$ reflects the proportion of generalized mentions whose annotation links share the same source domains as the recommended references. These two are assigned different weights and combined to compute the TrustworthyBoost $\in[1,1+\eta]$ as:
$\text{TBoost}=1+\eta\cdot\bigg[\ \theta\cdot\underbrace{\left(\frac{\text{match}_{\text{full}}}{S}\right)}_{\text{Rate}_{\text{full\,hit}}}+\kappa\cdot\underbrace{\left(\frac{\text{match}_{\text{host}}}{T+1}\right)}_{\text{Rate}_{\text{host\,hit}}}\ \bigg]$
where $\text{match}_{\text{full}}$ indicates the number that have been exactly matched in TSLs; $\text{match}_{\text{host}}$ refers to the number of annotations sharing the same hostname as the TSLs; $S$ and $T$ denote the sizes of the TSLs and the annotations respectively; $\theta$ and $\kappa$ represent the weights for full and hostname matches respectively, with $\theta+\kappa=1$. The coefficient $\eta$ controls the magnitude of the confidence boost, thereby preventing excessive inflation due to high confidence and avoiding complete nullification when confidence is low. This mechanism preserves scoring stability while enhancing evaluative sensitivity to verifiability and source reliability, thereby improving the responsiveness to external evidence signals.
@@ 4.4 Integrated Scoring Framework
The integrated metric $\in[0,120]$ adopts a multiplicative weighting model to enable multidimensional evaluation of report-style generation tasks.
IntegratedScore $=\text{Quality}\times(1-\text{SDrift})\times\text{TBoost}$
Here, Quality assesses the structural integrity and content quality of the report, SemanticDrift reflects the degree of thematic deviation and is transformed into a positive factor via $1-\text{SemanticDrift}$, and TrustworthyBoost enhances the credibility weight based on authoritative link coverage. This design logic penalizes semantic drift and rewards external support, producing a normalized score on a 100-point scale.
@@
26:  $\overline{\text{InteScore}}\leftarrow$ average of InteScore over entries
27:  $\overline{\text{ContriPerToken}}\leftarrow$ average of ContriPerToken over entries
Table 1: Results of the main experiment, with the highest score in bold and the second-highest underlined in each column.
Models Quality $\uparrow$ 11-SDrift $\uparrow$ TBoost $\uparrow$ InteScore $\uparrow$ Usage C/Token $\downarrow$
@@
4.5 Additional Metrics
To evaluate cost-effectiveness under limited resources, the Contribution/Token metric evaluates model efficiency based on actual token expenditure in reasoning and generation, effectively measuring the information density per token. When long texts maintain high information density, efficiency scores can be high; they are truly lowered only when the text contains substantial redundancy or off-topic content, which is precisely what we aim to detect. This metric is designed to encourage models to produce efficient outputs.
5 Experiments
5.1 Overview
We evaluated a total of thirteen models under Dr. Bench, including five DRAs (o3-deep-research-2025-06-
26
, qwen-deep-research, sonar-deep-resea-
rch
, grok-4-0709-search, and o4-mini-deep-
research-2025-06-26
), one advanced agent (kimi-
k2-0905-preview
), and seven web-search-tool-enhanced reasoning models (gemini-2.5-pro, gpt-
5-2025-08-07
, gpt-4o-search-preview-2025- 03-11, gpt-4.1-2025-04-14, claude-opus-4-
1-20250805
, claude-sonnet-4- 20250514, and claude-3-7-sonnet-20250219).
To eliminate the randomness, all models with adjustable temperature were evaluated at $\text{temperature}=0.0$. For non-DRA models without embedded annotations, reports were merged with annotations during evaluation. gpt-4o-20-
24-11-20
independently scored each rubric to evaluate semantic quality. Topical focus was assessed on pure report text without annotations to avoid interference from annotation titles. Related prompts are provided in Appendix D. Retrieval Trustworthiness was computed from pure annotations, excluding parameters, anchors, and duplicates to ensure fair matching. Manual verification was randomly conducted on approximately 35% of the scores judged by LLMs, yielding a 99.3% agreement with human evaluations.
5.2 Leaderboard
Table 1 and Appendix E presents the leaderboard results, with models ranked in descending order of IntegratedScore. In terms of IntegratedScore, Qwen ranked first, demonstrating strong performance across all dimensions. Sonar followed closely, achieving the highest score in topical focus. Notably, Kimi-K2, a Mixture-of-Experts architecture agent with 1T parameters, attained the highest score in the quality, outperforming all DRAs. For topical focus, Sonar, Qwen, and o3 formed the leading cluster. GPT-5 achieved the highest score in citation reliability, indicating strong external support alignment. In terms of resource consumption, o3 and o4-mini averaged 23K and 18K tokens per report respectively, placing them lowest in contribution efficiency.
Overall, DRAs exhibited stable semantic quality and control in structured report generation, while web-search-tool-augmented models showed promise in external support and efficiency. These results validate the Dr. Bench’s ability to differentiate model capabilities across multiple dimensions, providing a robust empirical foundation for future optimization of DRAs.
Figure 3: The proportion of quality share across models.
Figures 3 and 10 summarize the role of Quality scores and cross-domain performance in the overall IntegratedScore. While Kimi-K2 achieves a strong Quality score, its relative weaknesses in credibility and attention reduce the impact of this advantage, whereas models such as Qwen, Sonar, and o3 exhibit a more balanced interplay between Quality and multiplicative factors, yielding stronger integrated performance. Across domains, all top models perform notably well in domain 03 (Sports & Competitions) and domain 10 (Health & Medicine), while maintaining relatively balanced results elsewhere. In Quality, Kimi-K2 leads, followed by Qwen and Sonar; however, in SemanticDrift and TrustworthyBoost, Kimi-K2 shows clear disadvantages, whereas the others remain comparatively consistent. Detailed analysis is provided in Appendix E.
@@ 5.3 Supplementary Dimensions
Table 2: Average inference and retrieval times on OpenAI models.
@@
As shown in Table 2, GPT-4.1 showed minimal retrieval activity, with an average of only 0.39 times. In contrast, both o4-mini-dr and o3-dr displayed intensive inference and retrieval patterns, suggesting more complex reasoning chains and information acquisition strategies. In terms of retrieval efficiency, o3-dr slightly exceeded in total retrieved links and annotations, while o4-mini-dr achieved slightly lower RetrievalIndex, reflecting stronger filtering and citation precision, as shown in Table 3.
Table 3: RetrievalIndex of o3-dr and o4-mini-dr.
Models $\text{{num}}_{\text{{retr}}}$ $\text{{num}}_{\text{{anno}}}$ RIndex $\downarrow$
@@ 7 Conclusions
In this paper, we present the Dr. Bench and the multidimensional evaluation framework that systematically assesses the performance and capabilities of DRAs. By leveraging challenging queries across diverse thematic domains and high-quality reference bundles, our framework enables rigorous evaluation of report-style outputs along axes of semantic quality, topical focus, and retrieval trustworthiness. Empirical results show that contemporary DRAs substantially outperform conventional tool-augmented models in complex task scenarios, while also exposing key limitations and trade-offs. These insights elucidate current challenges in DRA design and lay the groundwork for advancing DRAs as efficient, stable, and interpretable intelligent agents.
Appendix A Taxonomy of Domains
@@ 00 Unclassified Cannot be categorized 3
Appendix B General-Report Rubrics
@@
Appendix C Examples of Entries
To illustrate the precision and rigor of our benchmark, we present four representative entries: ID 04216, 07001, 08004, and 10236. These entries span distinct domains and query types and serve as concrete examples of structural completeness, rubric coverage, and citation fidelity.
Figure 6: Example ID 07001 of benchmark entries.
Figure 7: Example ID 04216 of benchmark entries.
@@ Appendix C Examples of Entries
Figure 9: Example ID 10236 of benchmark entries.
Appendix D Scoring Prompts
@@ 2. A keyword to evaluate Your task is to: - Read the report carefully
Appendix E Supplementary Experimental Observations
E.1 Quality Share Proportion
@@
E.2 Evaluation Across Domains
@@ D M QWE SON O3D KIM GRO GEM O4D GT5 G4O G41 OPU SO4 S37
Appendix F Comparison with Existing Benchmarks
Table 6 presents a comparison of our benchmark with existing ones across key dimensions. Task Type specifies whether the task is open-ended or closed and whether the target response takes the form of a short string or a long report. Evaluation Dimensions indicate which metrics are applied and which aspects of model behavior are examined. Construction method describes whether the benchmark is primarily built through LLM-based generation or through human annotation. Scale refers to the size of the dataset in terms of the number of instances. Target System specifies whether the benchmark is designed for LLMs equipped with web-search tools or for DRAs. Finally, Evaluation Criteria clarify whether performance is judged by exact matching, LLM-based criteria, or expert rubric assessment.
Table 6: Comparison of Benchmarks for Tool-Augmented LLMs
Benchmarks Task Type EvalDim Constru Scale System EvalCriteria
WebWalker  (Wu et al., 2025) Closed/String Accuracy LLMs 680 LLMs Match
BrowseComp-Plus (Chen et al., 2025b) Closed/String Accuracy LLMs 830 DRAs Match
GAIA (Mialon et al., 2023) Closed/String Accuracy Human 466 LLMs Match
BrowseComp (Wei et al., 2025) Closed/String Accuracy Human 1266 LLMs Match
WideSearch (Wong et al., 2025) Closed/String F1 Score Human 200 LLMs Match
Deep Research Bench (Bosse et al., 2025) Closed/String Recall, F1 Score Human 89 LLMs Match
ResearchQA (Du et al., 2025, Concurrent Work) Open/Report Quality LLMs 21K DRAs LLMs Criteria
ReportBench (Li et al., 2025, Concurrent Work) Open/Report Quality LLMs 678 DRAs LLMs Criteria
DeepResearch Arena (Wan et al., 2025, Concurrent Work) Open/Report Quality LLMs 10K DRAs LLMs Criteria
DeepResearch Bench (Du et al., 2025, Concurrent Work) Open/Report Quality Human 100 DRAs LLMs Criteria
Dr. Bench (ours) Open/Report Quality, Semantics, Retrieval Human 214 DRAs Experts Rubrics, Keywords, Links
Dr. Bench is the only benchmark that combines quality, semantic adequacy, and retrieval credibility in evaluating report-style outputs. Unlike large-scale auto-generated datasets, it is human-authored with expert rubrics, anchors, and trustworthy links, ensuring transparency and reproducibility. Its 214 carefully curated tasks balance coverage with precision, making it uniquely suited for systematic and fine-grained assessment of Deep Research Agents. Our evaluation dimensions and methodology differ substantially from prior work, with more qualitative and fine-grained details provided in Section 2.2.
Appendix G Core Dimensions of QSRs
Within the evaluation framework for structured long-text tasks, a critical criterion for assessing the effectiveness and representativeness of QSRs is whether their design demonstrates discriminative capacity in judging task completeness and aligns with human user expectations. QSRs constitute a dimension-based framework tailored to specific queries, intended to measure the completeness of reports. The design rationale of QSRs considers, though is not limited to, the core dimensions outlined in Table 7.
Table 7: Core dimensions with descriptions of QSRs
Dimensions Descriptions
Information Coverage Whether the report fully covers key facts, events, clauses, persons, institutions, or geographic units required by the query, ensuring no essential information is missing.
Technical / Mechanism Explanation Whether the report accurately explains mechanisms, protocols, causal chains, policy tools, or theoretical frameworks involved in the query, reflecting mastery of core logic.
Structural Expression Whether the report generates structured forms such as tables, timelines, maps, matrices, or lists to support organization and reproducibility of the task.
Semantic Precision Whether the report defines key terms, distinguishes concepts, annotates units and time, and avoids ambiguity, ensuring clarity and consistency of expression.
Source Verification Whether the report cites authoritative texts, identifiers, database versions, retrieval times, links, or snapshots to support traceability and credibility of claims.
Evidence Organization Whether the report employs multi-source cross-validation, independent source pairing, counterfactual baselines, placebo tests, or robustness checks to enhance reliability.
Heterogeneity and Comparative Analysis Whether the report identifies and analyzes differences across regions, groups, time periods, or institutions, supporting comparative and explanatory power.
Methodological Transparency Whether the report clearly explains analytical methods, identification strategies, variable definitions, and data processing steps, ensuring consistency and reproducibility.
Social Impact and Distribution Effects Whether the report analyzes distributional impacts of policies or events on different groups, communities, industries, or ecosystems, reflecting social dimensions.
Historical Evolution and Temporal Logic Whether the report establishes event chains, policy evolution paths, time windows, or lag mechanisms, supporting historical and dynamic analysis.
Interdisciplinary Integration Whether the report integrates perspectives from multiple disciplines (e.g., law, economics, technology, society, philosophy) to support complexity and explanatory depth.